Disk Structure

July 1, 1997

The layout of my hard disks, as of 1/7/97 is as follows.

Disk /dev/hda: 64 heads, 63 sectors, 1023 cylinders
Units = cylinders of 4032 * 512 bytes
   Device Boot   Begin    Start      End   Blocks   Id  System
/dev/hda1   *        1        1      256   516064+   6  DOS 16-bit >=32M
/dev/hda2          257      257      512   516096    5  Extended
/dev/hda3          513      513      545    66528   82  Linux swap
/dev/hda4          546      546     1023   963648   83  Linux native
/dev/hda5          257      257      512   516064+   6  DOS 16-bit >=32M

Disk /dev/hdd: 16 heads, 63 sectors, 4095 cylinders
Units = cylinders of 1008 * 512 bytes
   Device Boot   Begin    Start      End   Blocks   Id  System
/dev/hdd1            1        1     1220   614848+   6  DOS 16-bit >=32M
/dev/hdd2         2868     3277     4088   409248    5  Extended
/dev/hdd3         1024     1221     2044   415296   83  Linux native
/dev/hdd4         1024     2045     2867   414792   83  Linux native
/dev/hdd5         2868     3277     4088   409216+   6  DOS 16-bit >=32M

Note, the second hard drive has the following anomylies because the number of cylinders is 4095¿1024.

Partition 1 does not end on cylinder boundary:
     phys=(304, 63, 63) should be (304, 15, 63)
Partition 2 does not end on cylinder boundary:
     phys=(1021, 63, 63) should be (1021, 15, 63)

My fstab file is as follows.

/dev/hda3        swap          swap       defaults 1 1
/dev/hda4        /             ext2       defaults 1 1
/dev/hda1        /win_c        msdos      defaults 1 1
/dev/hda5        /win_e        msdos      defaults 1 1
/dev/hdd1        /win_d        msdos      defaults 1 1
/dev/hdd5        /win_f        msdos      defaults 1 1
/dev/hdd3        /sundance     ext2       defaults 1 1
/dev/hdd4        /drizzt       ext2       defaults 1 1
none             /proc         proc       defaults 1 1